OpenStudy (anonymous):

WILL MEDAL Solve x3 = 64/27.

OpenStudy (anonymous):

64/81=x

OpenStudy (anonymous):

@firedupme did u get it

OpenStudy (anonymous):

??!

OpenStudy (anonymous):

hold on

OpenStudy (mrnood):

\[x ^{3} =\frac{ 64 }{ 27 }\] so take cube root to find x \[\sqrt[3]{x ^{3}} = x = \sqrt[3]{\frac{ 64 }{ 27}} = \frac{ \sqrt[3]{64} }{\sqrt[3]{27} }\]

OpenStudy (anonymous):

@MrNood is correct. He got it to it before me :P good job.

OpenStudy (mrnood):

@moha_10 you have misread the question - it is x^3 - NOT 3x

OpenStudy (anonymous):

im pretty sure my answer is wrong

OpenStudy (anonymous):

im terrible at these questions

OpenStudy (anonymous):

alright is my mistake sorry @MrNood

OpenStudy (anonymous):

the answers are ±8 / 3 8 / 3 ±4 / 3 4 / 3

OpenStudy (anonymous):

@MrNood im really bad at these so just bare with me.

OpenStudy (mrnood):

I wrote all but the last step above the answer is \[x= \sqrt[3]{64}\div \sqrt[3]{27}\] it turns out that those 2 numbers are VERY easy to get the cube root use your calculator - or just use some common sense what number is cubed to make 64? what number is cubed to make 27?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

8/3

OpenStudy (anonymous):

what about the ±?

OpenStudy (mrnood):

8 cubed = 8x8x8 = 492

OpenStudy (anonymous):

so the answer is 8/3

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

PLZ

OpenStudy (mrnood):

I just showed you that 8 cubed is NOT 64 so therefore cuberoot of 64 is not 8 so 8 is not in the answer

OpenStudy (anonymous):

so it would have to be 4/3. i just don't get the ±

OpenStudy (anonymous):

so is it ±4/3 or just 4/3 ?

OpenStudy (mrnood):

as an example \[\sqrt{4} =\pm 2\] BECAUSE -2 x -2 = 4 AND 2 X 2 = 4 is this true for CUBE root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so ±4/3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Nope that waws wrong ,,,, fun